4x + 7y = 8.77
8x + 14y = 15.8
I ued substitution and got x=1.2175 and y=39 over 70. Obviously I did something wrong.

Answer by mananth(16946)

4 x + 7 y = 8.77 .............1
8 x + 14 y = 15.8 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-8 x -14 y = -17.54
8 x 14 y = 15.8
Add the two equations
0 x = -1.74
which is absurd
Hence there is no solution to the system.
The lines are parallel